OpenStudy (anonymous):

What is the 20th term of the arithmetic sequence 4, -2, -8, ... ? :p -110 -116 -122 -128

OpenStudy (zepp):

First, what is your common ratio here?

OpenStudy (anonymous):

2?

OpenStudy (zepp):

2? Are you sure? :P

OpenStudy (zepp):

What did you we do to 4 so it becomes a -2?

OpenStudy (anonymous):

if this is arithmetic, shouldn't it be common difference?

OpenStudy (zepp):

Oh, right :( I just like the term 'common ratio' too much :U

OpenStudy (anonymous):

im dont really understand what to do next sorry?

OpenStudy (zepp):

4 - 6 --> -2 Our common difference would be -6 here.

OpenStudy (anonymous):

*i

OpenStudy (anonymous):

ohh ok

OpenStudy (zepp):

The use the formula \(x_n = a+(n-1)d\) Where \(x_n\) is what you are looking for; a the first term; n the term you are looking for; d the common difference. We want the 20th, let n be 20 \(x_{20} = a+(20-1)d\) And plug in everything else \(x_{20} = 4+(20-1)(-6)\)